OpenStudy (anonymous):

5+(x-12)=18 What is the solution to the equation?

OpenStudy (richgall____):

x=25

OpenStudy (richgall____):

25-12=13+5=18

OpenStudy (anonymous):

THANK U

OpenStudy (richgall____):

Welcome.